Given the demand function P = 20 – 5Q, find the price elasticity of demand when price of the commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point
1
Expert's answer
2021-08-18T15:50:29-0400

Price elasticity of demand

"=\frac{Q_2-Q_1}{\frac {Q_2+Q_1}{2}} \times100"

"=\frac {3-3.8}{\frac {3+3.8}{2}}\times100"

"=-23.53"

"\frac {P_2-P_1}{\frac {P_2+P_1}{2}}\times 100"

"\frac {5-1}{\frac {5+1}{2}}\times100 =133.33"

"\therefore" Price elasticity of demand

"=\frac {-23.53}{133.33}"

"=0.18"

The price elasticity of demand is inelastic.